# Flavored QCD axion and Modular invariance

### v1

@@ Abstract
A four-dimensional effective model with $G_{\rm SM}\times SL(2,\mathbb{Z})\times U(1)_{X}$ is proposed in string-derived supergravity framework, where $G_{\rm SM}$ is the Standard Model (SM) gauge group and $U(1)_{X}$ is gauged. We show $SL(2,\mathbb{Z})$- and $U(1)_{X}$-mixed anomalies should vanish. Anomalies induced by Kähler transformations match those from gaugino chiral rotations. When SM fermions transform nontrivially under $SL(2,\mathbb{Z})$, and with vanishing gaugino contributions, the anomaly-free conditions are powerful enough to determine the quark and lepton flavor structures, set scales for $U(1)_{X}$ breaking, and ensure the strong CP phase remains unmodified. While the Green-Schwarz coefficient $\delta^{\rm GS}_{X}$ is generically non-zero, vanishing $U(1)_{X}$ anomalies cause gauge boson decoupling and $\delta^{\rm GS}_{X}\rightarrow 0$, yielding a massless global $U(1)_{X}$ without a Nambu-Goldstone mode. We show that the modulus vacuum expectation value stabilizes near $\langle\tau\rangle\approx i$, where exact $SL(2,\mathbb{Z})$ ($T$-duality) is spontaneously broken, removing residual modular symmetry. The framework predicts seesaw-generated neutrino masses and flavored axion properties, with all Yukawa coefficients constrained to unit-magnitude complex numbers. Our model reproduces current quark and lepton data, predicts an axion mass $m_{a}=1.82\times 10^{-4}$ eV and photon coupling $|g_{a\gamma\gamma}|=1.75\times 10^{-14}\,{\rm GeV}^{-1}$, and unlike the ordinary case, suppresses flavor-violating axion couplings to $s,d$ quarks and $\mu,e$ leptons to $\mathcal{O}(\lambda^{4})$ (with $\lambda$ the Cabibbo angle). It also yields normal neutrino mass hierarchy consistent with oscillation data, $0\nu\beta\beta$-decay rate, and cosmological and astrophysical measurements.
@@ III.2 Modular-invariant Yukawa superpotential for lepton
$U(1)_{X}$ $\frac{1}{2}$ $\frac{1}{2}$ $\frac{1}{2}$ $\frac{1}{2}+23$ $\frac{1}{2}+11$ $\frac{1}{2}+7$ $\frac{1}{2}$ $\frac{1}{2}$ $\frac{1}{2}$
@@
$W_{\ell}$ $=$ $\Big[\alpha_{\tau}\Big(\frac{\tilde{\chi}}{\Lambda}\Big)^{7}\tau^{c}L_{\tau}+\alpha_{\mu}\Big(\frac{\tilde{\chi}}{\Lambda}\Big)^{11}\mu^{c}L_{\mu}+\alpha_{e}\Big(\frac{\tilde{\chi}}{\Lambda}\Big)^{23}e^{c}L_{e}$ (55)
$+\alpha_{\tau 2}\Big(\frac{\tilde{\chi}}{\Lambda}\Big)^{7}Y^{(8)}_{\bf 1}\tau^{c}L_{\mu}+\alpha_{\tau 1}\Big(\frac{\tilde{\chi}}{\Lambda}\Big)^{7}Y^{(4)}_{\bf 1}\tau^{c}L_{e}+\alpha_{e2}\Big(\frac{\tilde{\chi}}{\Lambda}\Big)^{23}Y^{(4)}_{\bf 1}e^{c}L_{\mu}\Big]H_{d}+.$..\,.
@@
$W_{\nu}$ $=$ $\Big[\beta_{1e}Y^{(24)}_{\bf 1}N^{c}_{1}L_{e}+\beta_{1\mu}Y^{(28)}_{\bf 1}N^{c}_{1}L_{\mu}+\beta_{1\tau}Y^{(20)}_{\bf 1}N^{c}_{1}L_{\tau}$ (56)
@@
$E$ $=$ $\delta^{W}_{X}+\delta^{Y}_{X}=\frac{494}{3}\,$, (58)
where the anomaly coefficients $\delta^{W}_{X}$, $\delta^{Y}_{X}$ are defined in Eq.(31). The physical quantities of QCD axion, such as axion mass $m_{a}$ and axion-photon coupling $g_{a\gamma\gamma}$, depend on the ratio of electromagnetic anomaly coefficient $E$ to the color anomaly coefficient $\delta^{G}_{X}$ (see Fig.2). This example model has the specific ratio $E/\delta^{G}_{X}=494/201$ with the $U(1)_{X}$ breaking scale fixed at $f_{A}=2\times 10^{12}$ GeV (as derived in Eq.(97)). This distinctive prediction makes the model testable in future axion search experiments.
@@ IV Quark, Lepton, and Flavored-QCD axion
$+$ $\overline{\ell_{R}}\,{\cal M}_{\ell}\,\ell_{L}+\frac{1}{2}\begin{pmatrix}\overline{\nu^{c}_{L}}&\overline{N_{R}}\end{pmatrix}\begin{pmatrix}0&m^{T}_{D}\\ m_{D}&~~~e^{i\frac{A_{X}}{f_{A}}}\,M_{R}\end{pmatrix}\begin{pmatrix}\nu_{L}\\ N^{c}_{R}\end{pmatrix}+\frac{g}{\sqrt{2}}W^{-}_{\mu}\overline{\ell_{L}}\gamma^{\mu}\,\nu_{L}+\text{h.c.}\,$,
@@
${\cal M}_{\ell}$ $=$ $C^{\ell}_{R}{\left(\begin{array}[]{ccc}\alpha_{e}\,\Delta^{23}_{\chi}&\alpha_{e2}Y^{(4)}_{\bf 1}\Delta^{23}_{\chi}(2{\rm Im}\,\tau)^{-2}&0\\ 0&\alpha_{\mu}\,\Delta^{11}_{\chi}&0\\ \alpha_{\tau 1}Y^{(4)}_{\bf 1}\Delta^{7}_{\chi}(2{\rm Im}\,\tau)^{-2}&\alpha_{\tau 2}Y^{(8)}_{\bf 1}\Delta^{7}_{\chi}(2{\rm Im}\,\tau)^{-4}&\alpha_{\tau}\,\Delta^{7}_{\chi}\end{array}\right)}\,v_{d}\,$, (73)
$M_{R}$ $=$ $(2{\rm Im}\,\tau)^{-6}{\left(\begin{array}[]{ccc}\gamma_{11}\,Y^{(12)}_{{\bf 1}}&\gamma_{12}\,Y^{(12)}_{{\bf 1}}&\gamma_{13}\,Y^{(12)}_{{\bf 1}}\\ \gamma_{12}\,Y^{(12)}_{{\bf 1}}&\gamma_{22}\,Y^{(12)}_{{\bf 1}}&\gamma_{23}\,Y^{(12)}_{{\bf 1}}\\ \gamma_{13}\,Y^{(12)}_{{\bf 1}}&\gamma_{23}\,Y^{(12)}_{{\bf 1}}&\gamma_{33}\,Y^{(12)}_{{\bf 1}}\end{array}\right)}\langle\chi\rangle\,$, (77)
$m_{D}={\left(\begin{array}[]{ccc}\beta_{1e}\,Y^{(24)}_{{\bf 1}}(2{\rm Im}\,\tau)^{-12}&\beta_{1\mu}\,Y^{(28)}_{\bf 1}(2{\rm Im}\,\tau)^{-14}&\beta_{1\tau}\,Y^{(20)}_{\bf 1}(2{\rm Im}\,\tau)^{-10}\\ \beta_{2e}\,Y^{(24)}_{{\bf 1}}(2{\rm Im}\,\tau)^{-12}&\beta_{2\mu}\,Y^{(28)}_{{\bf 1}}(2{\rm Im}\,\tau)^{-14}&\beta_{2\tau}\,Y^{(20)}_{{\bf 1}}(2{\rm Im}\,\tau)^{-10}\\ \beta_{3e}\,Y^{(24)}_{\bf 1}(2{\rm Im}\,\tau)^{-12}&\beta_{3\mu}\,Y^{(28)}_{\bf 1}(2{\rm Im}\,\tau)^{-14}&\beta_{3\tau}Y^{(20)}_{{\bf 1}}(2{\rm Im}\,\tau)^{-10}\end{array}\right)}\,v_{u}\,$, (81)
@@
$C^{\ell}_{R}={\rm diag}(e^{-i23\frac{A_{X}}{f_{A}}},e^{-i11\frac{A_{X}}{f_{A}}},e^{-i7\frac{A_{X}}{f_{A}}})\,.$ (82)
@@ IV.1 Quark and charged-lepton masses, mixing, and QCD axion interactions
$m_{e}\simeq|\alpha_{e}|\,\Delta^{23}_{\chi}\,v_{d}\,,\qquad m_{\mu}\simeq|\alpha_{\mu}|\,\Delta^{11}_{\chi}\,v_{d}\,,\qquad m_{\tau}\simeq|\alpha_{\tau}|\,\Delta^{7}_{\chi}\,v_{d}\,.$ (85)
@@
[…]
@@
$-{\cal L}^{a\ell}\simeq$ - $\frac{\partial_{\mu}a}{2f_{A}}\Big\{23\,\bar{e}\gamma^{\mu}\gamma_{5}e+(11-4\theta^{r2}_{23})\,\bar{\mu}\gamma^{\mu}\gamma_{5}\mu+(7+4\theta^{r2}_{23})\,\bar{\tau}\gamma^{\mu}\gamma_{5}\tau\Big\}$ (87)
- $\frac{\partial_{\mu}a}{2f_{A}}\Big\{4\theta^{r}_{23}e^{i(\alpha^{\ell}_{2}-\alpha^{\ell}_{3})}\,\bar{\mu}\gamma^{\mu}(1+\gamma_{5})\tau+{\rm h.c.}\Big\}+\sum_{\ell=e,\mu,\tau}\big(m_{\ell}\bar{\ell}\ell-\bar{\ell}i\!\!\not\!\partial\,\ell\big)\,.$
with mixing matrix governed by $V^{\ell}_{R}$ in Eq.(138). The flavored-QCD axion $a$ is produced by flavor-changing neutral Yukawa interactions in Eqs.(86) and (87), which leads to induced rare flavor-changing processes such as $b\rightarrow s+a$, $b\rightarrow d+a$CLEO:2001acz , and $\tau\rightarrow\mu+a$BaBar:2009hkt ; ARGUS:1995bjh . In our flavored-axion model, the bounds on $f_{A}$ arise from $B^{\pm}\to K^{\pm}+a$ and/or $\tau\to\mu+a$ Bjorkeroth:2018dzu ; Calibbi:2016hwq ; delaVega:2021ugs , typically requiring $f_{A}\gtrsim 10^{5-6}$ GeV, rather than $K^{+}\rightarrow\pi^{+}+a$Wilczek:1982rv ; Bolton:1988af ; Artamonov:2008qb ; raredecay ; Berezhiani:1989fp . Flavor-violating couplings of the flavored QCD axion to $s,d$-quarks and $\mu,e$-leptons are suppressed to $\mathcal{O}(\lambda^{4})$, as shown in Eqs.(86) and (87). This suppression is a direct consequence of the $U(1)_{X}$ charge assignments  [^17] See $C^{d}_{L}$ in Eq.(69). and the following properties of the mixing matrices: the right-handed down-type quark matrix is approximately unitary up to $\mathcal{O}(\lambda^{4})$ (for $K^{+}\rightarrow\pi^{+}+a$), and the right-handed charged-lepton matrix has a 2-3 mixing angle of $\sim\mathcal{O}(\lambda^{2})$ (for $\mu\rightarrow e+a$), as detailed in Appendix E and Eq.(138).
The axion-electron coupling, given by $g_{aee}=23m_{e}/f_{A}$, is constrained by red giant branch (RGB) stars Viaux:2013lha :
$|g_{aee}|<4.3\times 10^{-13}\quad(95\%\,\text{CL})\quad\Leftrightarrow\quad\delta^{G}_{X}F_{a}\gtrsim 2.73\times 10^{10}\,\text{GeV}.$ (88)
@@
Figure 2: Plot of $|g_{a\gamma\gamma}|$ versus $m_{a}$ for Kim-Shifman-Vainshtein-Zakharov (KSVZ) KSVZ (black dashed line), Dine-Fischler-Srednicki-Zhitnitsky (DFSZ) DFSZ (blue dotted line), and the model (red solid line) in terms of $E/\delta^{G}_{X}=0,8/3$, and $494/201$, respectively. With $f_{A}=2\times 10^{12}$ GeV of Eq.(97) and $z=0.47$ our model predicts $m_{a}=1.82\times 10^{-4}$ eV and $|g_{a\gamma\gamma}|=1.75\times 10^{-14}\,{\rm GeV}^{-1}$ (black point).
@@
$f_{A}=2\times 10^{12}\,{\rm GeV}\,$, (92)
for $\langle\chi\rangle=10^{12}$ GeV in Eq.(97). The flavored-QCD axion mass and axion-photon coupling are predicted, as depicted in Fig.2, as
$m_{a}=1.82\times 10^{-4}\,{\rm eV}\,,\qquad|g_{a\gamma\gamma}|=1.75\times 10^{-14}\,{\rm GeV}^{-1}\,.$ (93)

### v2

@@ Abstract
A four-dimensional effective model with $G_{\rm SM}\times SL(2,\mathbb{Z})\times U(1)_{X}$ is proposed in string-derived supergravity framework, where $G_{\rm SM}$ is the Standard Model (SM) gauge group and $U(1)_{X}$ is gauged. We show $SL(2,\mathbb{Z})$- and $U(1)_{X}$-mixed anomalies should vanish. Anomalies induced by Kähler transformations match those from gaugino chiral rotations. When SM fermions transform nontrivially under $SL(2,\mathbb{Z})$, and with vanishing gaugino contributions, the anomaly-free conditions are powerful enough to determine the quark and lepton flavor structures, set scales for $U(1)_{X}$ breaking, and ensure the strong CP phase remains unmodified. While the Green-Schwarz coefficient $\delta^{\rm GS}_{X}$ is generically non-zero, vanishing $U(1)_{X}$ anomalies cause gauge boson decoupling and $\delta^{\rm GS}_{X}\rightarrow 0$, yielding a massless global $U(1)_{X}$ without a Nambu-Goldstone mode. We show that the modulus vacuum expectation value stabilizes near $\langle\tau\rangle\approx i$, where exact $SL(2,\mathbb{Z})$ ($T$-duality) is spontaneously broken, removing residual modular symmetry. The framework predicts seesaw-generated neutrino masses and flavored axion properties, with all Yukawa coefficients constrained to unit-magnitude complex numbers. Our model reproduces current quark and lepton data, predicts an axion mass $m_{a}\approx 0.9\times 10^{-2}$ eV and photon coupling $|g_{a\gamma\gamma}|\approx 1.7\times 10^{-13}\,{\rm GeV}^{-1}$, and unlike the ordinary case, suppresses flavor-violating axion couplings to $s,d$ quarks and $\mu,e$ leptons to $\mathcal{O}(\lambda^{4})$ (with $\lambda$ the Cabibbo angle). It also yields normal neutrino mass hierarchy consistent with oscillation data, $0\nu\beta\beta$-decay rate, and cosmological and astrophysical measurements.
@@ III.2 Modular-invariant Yukawa superpotential for lepton
$U(1)_{X}$ $\frac{1}{2}+4$ $\frac{1}{2}+4$ $\frac{1}{2}+4$ $\frac{1}{2}+20$ $\frac{1}{2}-15$ $\frac{1}{2}-8$ $\frac{1}{2}$ $\frac{1}{2}$ $\frac{1}{2}$
@@
$W_{\ell}$ $=$ $\Big[\alpha_{\tau}\Big(\frac{\chi}{\Lambda}\Big)^{4}\tau^{c}L_{\tau}+\alpha_{\mu}\Big(\frac{\chi}{\Lambda}\Big)^{11}\mu^{c}L_{\mu}+\alpha_{e}\Big(\frac{\tilde{\chi}}{\Lambda}\Big)^{24}e^{c}L_{e}$ (55)
$+\alpha_{\tau 2}\Big(\frac{\chi}{\Lambda}\Big)^{4}Y^{(8)}_{\bf 1}\tau^{c}L_{\mu}+\alpha_{\tau 1}\Big(\frac{\chi}{\Lambda}\Big)^{4}Y^{(4)}_{\bf 1}\tau^{c}L_{e}+\alpha_{e2}\Big(\frac{\tilde{\chi}}{\Lambda}\Big)^{24}Y^{(4)}_{\bf 1}e^{c}L_{\mu}\Big]H_{d}+.$..\,.
@@
$W_{\nu}$ $=$ $\Big(\frac{\tilde{\chi}}{\Lambda}\Big)^{4}\Big[\beta_{1e}Y^{(24)}_{\bf 1}N^{c}_{1}L_{e}+\beta_{1\mu}Y^{(28)}_{\bf 1}N^{c}_{1}L_{\mu}+\beta_{1\tau}Y^{(20)}_{\bf 1}N^{c}_{1}L_{\tau}$ (56)
@@
$E$ $=$ $\delta^{W}_{X}+\delta^{Y}_{X}=\frac{386}{3}\,$, (58)
where the anomaly coefficients $\delta^{W}_{X}$, $\delta^{Y}_{X}$ are defined in Eq.(31). The physical quantities of QCD axion, such as axion mass $m_{a}$ and axion-photon coupling $g_{a\gamma\gamma}$, depend on the ratio of electromagnetic anomaly coefficient $E$ to the color anomaly coefficient $\delta^{G}_{X}$ (see Fig.2). This example model has the specific ratio $E/\delta^{G}_{X}=386/201$ with the $U(1)_{X}$ breaking scale fixed at $f_{A}=4\times 10^{10}$ GeV (as derived in Eq.(97)). This distinctive prediction makes the model testable in future axion search experiments.
@@ IV Quark, Lepton, and Flavored-QCD axion
$+$ $\overline{\ell_{R}}\,{\cal M}_{\ell}\,\ell_{L}+\frac{1}{2}\begin{pmatrix}\overline{\nu^{c}_{L}}&\overline{N_{R}}\end{pmatrix}\begin{pmatrix}0&m^{T}_{D}\\ m_{D}&M_{R}\end{pmatrix}\begin{pmatrix}\nu_{L}\\ N^{c}_{R}\end{pmatrix}+\frac{g}{\sqrt{2}}W^{-}_{\mu}\overline{\ell_{L}}\gamma^{\mu}\,\nu_{L}+\text{h.c.}\,$,
@@
${\cal M}_{\ell}$ $=$ $C^{\ell}_{R}{\left(\begin{array}[]{ccc}\alpha_{e}\,\Delta^{24}_{\chi}&\alpha_{e2}Y^{(4)}_{\bf 1}\Delta^{24}_{\chi}(2{\rm Im}\,\tau)^{-2}&0\\ 0&\alpha_{\mu}\,\Delta^{11}_{\chi}&0\\ \alpha_{\tau 1}Y^{(4)}_{\bf 1}\Delta^{4}_{\chi}(2{\rm Im}\,\tau)^{-2}&\alpha_{\tau 2}Y^{(8)}_{\bf 1}\Delta^{4}_{\chi}(2{\rm Im}\,\tau)^{-4}&\alpha_{\tau}\,\Delta^{4}_{\chi}\end{array}\right)}\,v_{d}\,$, (73)
$M_{R}$ $=$ $e^{i\frac{A_{X}}{f_{A}}}\,(2{\rm Im}\,\tau)^{-6}{\left(\begin{array}[]{ccc}\gamma_{11}\,Y^{(12)}_{{\bf 1}}&\gamma_{12}\,Y^{(12)}_{{\bf 1}}&\gamma_{13}\,Y^{(12)}_{{\bf 1}}\\ \gamma_{12}\,Y^{(12)}_{{\bf 1}}&\gamma_{22}\,Y^{(12)}_{{\bf 1}}&\gamma_{23}\,Y^{(12)}_{{\bf 1}}\\ \gamma_{13}\,Y^{(12)}_{{\bf 1}}&\gamma_{23}\,Y^{(12)}_{{\bf 1}}&\gamma_{33}\,Y^{(12)}_{{\bf 1}}\end{array}\right)}\langle\chi\rangle\,$, (77)
$m_{D}=e^{-4i\frac{A_{X}}{f_{A}}}\,\Delta^{4}_{\chi}{\left(\begin{array}[]{ccc}\beta_{1e}\,Y^{(24)}_{{\bf 1}}(2{\rm Im}\,\tau)^{-12}&\beta_{1\mu}\,Y^{(28)}_{\bf 1}(2{\rm Im}\,\tau)^{-14}&\beta_{1\tau}\,Y^{(20)}_{\bf 1}(2{\rm Im}\,\tau)^{-10}\\ \beta_{2e}\,Y^{(24)}_{{\bf 1}}(2{\rm Im}\,\tau)^{-12}&\beta_{2\mu}\,Y^{(28)}_{{\bf 1}}(2{\rm Im}\,\tau)^{-14}&\beta_{2\tau}\,Y^{(20)}_{{\bf 1}}(2{\rm Im}\,\tau)^{-10}\\ \beta_{3e}\,Y^{(24)}_{\bf 1}(2{\rm Im}\,\tau)^{-12}&\beta_{3\mu}\,Y^{(28)}_{\bf 1}(2{\rm Im}\,\tau)^{-14}&\beta_{3\tau}Y^{(20)}_{{\bf 1}}(2{\rm Im}\,\tau)^{-10}\end{array}\right)}\,v_{u}\,$, (81)
@@
$C^{\ell}_{R}={\rm diag}(e^{-i24\frac{A_{X}}{f_{A}}},e^{i11\frac{A_{X}}{f_{A}}},e^{i4\frac{A_{X}}{f_{A}}})\,.$ (82)
@@ IV.1 Quark and charged-lepton masses, mixing, and QCD axion interactions
$m_{e}\simeq|\alpha_{e}|\,\Delta^{24}_{\chi}\,v_{d}\,,\qquad m_{\mu}\simeq|\alpha_{\mu}|\,\Delta^{11}_{\chi}\,v_{d}\,,\qquad m_{\tau}\simeq|\alpha_{\tau}|\,\Delta^{4}_{\chi}\,v_{d}\,.$ (85)
@@
After diagonalizing the quark mass matrices of Eqs.(64,68), the leading-order flavored-QCD axion interactions with quarks (up to $\mathcal{O}(\lambda^{4})$) are  [^16] The left-handed quark and the right-handed charged-lepton mixing matrices, $V^{d(u)}_{L}$ and $V^{\ell}_{R}$, are determined in Appendix E. These matrices govern the flavored-QCD axion couplings to down-type quarks, up-type quarks, and charged leptons, respectively.
@@
$-{\cal L}^{a\ell}\simeq$ - $\frac{\partial_{\mu}a}{2f_{A}}\Big\{24\,\bar{e}\gamma^{\mu}\gamma_{5}e-11\,\bar{\mu}\gamma^{\mu}\gamma_{5}\mu-4\,\bar{\tau}\gamma^{\mu}\gamma_{5}\tau\Big\}+\sum_{\ell=e,\mu,\tau}\big(m_{\ell}\bar{\ell}\ell-\bar{\ell}i\!\!\not\!\partial\,\ell\big)\,.$ (87)
The flavored-QCD axion $a$ is produced by flavor-changing neutral Yukawa interactions in Eqs.(86) and (87), which leads to induced rare flavor-changing processes such as $b\rightarrow s+a$, $b\rightarrow d+a$CLEO:2001acz , and $\tau\rightarrow\mu+a$BaBar:2009hkt ; ARGUS:1995bjh . In our flavored-axion model, the bounds on $f_{A}$ arise from $B^{\pm}\to K^{\pm}+a$Bjorkeroth:2018dzu ; Calibbi:2016hwq ; delaVega:2021ugs , typically requiring $f_{A}\gtrsim 10^{5-6}$ GeV, rather than $K^{+}\rightarrow\pi^{+}+a$Wilczek:1982rv ; Bolton:1988af ; Artamonov:2008qb ; raredecay ; Berezhiani:1989fp . Flavor-violating couplings of the flavored QCD axion to $s,d$-quarks and $\mu,e$-leptons are suppressed to $\mathcal{O}(\lambda^{4})$, as shown in Eqs.(86) and (87). This suppression is a direct consequence of the $U(1)_{X}$ charge assignments  [^17] See $C^{d}_{L}$ in Eq.(69). and the following properties of the mixing matrices: the right-handed down-type quark matrix is approximately unitary up to $\mathcal{O}(\lambda^{4})$ (for $K^{+}\rightarrow\pi^{+}+a$), and the right-handed charged-lepton matrix is approximately the identity up to $\mathcal{O}(\lambda^{4})$ (for $\mu\rightarrow e+a$), as detailed in Appendix E.
The axion-electron coupling, given by $g_{aee}=24m_{e}/f_{A}$, is constrained by red giant branch (RGB) stars Viaux:2013lha :
$|g_{aee}|<4.3\times 10^{-13}\quad(95\%\,\text{CL})\quad\Leftrightarrow\quad\delta^{G}_{X}F_{a}\gtrsim 2.85\times 10^{10}\,\text{GeV}.$ (88)
@@
Figure 2: Plot of $|g_{a\gamma\gamma}|$ versus $m_{a}$ for Kim-Shifman-Vainshtein-Zakharov (KSVZ) KSVZ (black dashed line), Dine-Fischler-Srednicki-Zhitnitsky (DFSZ) DFSZ (blue dotted line), and the model (red solid line) in terms of $E/\delta^{G}_{X}=0,8/3$, and $386/201$, respectively. With $f_{A}=4\times 10^{10}$ GeV of Eq.(97) and $z=0.47$ our model predicts $m_{a}=9.12\times 10^{-3}$ eV and $|g_{a\gamma\gamma}|=1.69\times 10^{-13}\,{\rm GeV}^{-1}$ (black point).
@@
$f_{A}=4\times 10^{10}\,{\rm GeV}\,$, (92)
for $\langle\chi\rangle=2\times 10^{10}$ GeV in Eq.(97). The flavored-QCD axion mass and axion-photon coupling are predicted, as depicted in Fig.2, as
$m_{a}=9.12\times 10^{-3}\,{\rm eV}\,,\qquad|g_{a\gamma\gamma}|=1.69\times 10^{-13}\,{\rm GeV}^{-1}\,.$ (93)